OpenStudy (anonymous):

evaluate each expression if a=4 b=-10 and c=-4 3a+4b+6c

OpenStudy (anonymous):

3*4+4*(-10)+6*(-4)=12-40-24=-52

OpenStudy (anonymous):

please explain further don,t understand